OpenStudy (anonymous):

There are a total of 15 balls in a bag. The bag contains 3 balls of each of the following colours: red, green, blue, yellow and orange. For $5.00, you can choose 3 balls (You cannot put the ball back after choosing it). If you pick 3 balls and they all match in colour, you win the grand prize of 100 dollars. What is the probability of choosing 3 matching balls? I did: (3C3 x 3C0 x 3C0 x 3C0 x 3C0/ 15C3) x 5C1 = (1/455) x 5C1 = 1/91 What did I do wrong?

OpenStudy (ybarrap):

Are you trying to find the probability of winning?

OpenStudy (anonymous):

yes, of choosing 3 balls that match in color

OpenStudy (ybarrap):

So you have 5 colors. You pick a ball. It's a particular color. There are 14 balls, left. But only 2 of the 14 will result in a win. You pick a ball on the second draw of the same color as the 1st with probability: 2/14 You pick a ball on the third draw of the same color as the 1st with probability 1/13 So the probability that the 2nd and 3rd draws are of the same color as the 1st is: 2/14 * 1/13 = 2/182=1/91

OpenStudy (anonymous):

I get it, thank you!

OpenStudy (ybarrap):

you're welcome